in a family of 4 children. what is the probability that one boy and three girls? use P=a^sb^t)n!/(s!t!)
Answer #1

No.of children n = 4

Assuming probability of having a boy= probability of having a girl=1/2

P = 4C3*(1/2)^3*(1/2)

= 4/16

= 1/4

= 0.25